# NightHaze: Nighttime Image Dehazing via Self-Prior Learning

### v1

@@ 4.1 Implementation Details
For a haze image of size 256 $\times$ 256, we first sample overlapping regions within the input haze image. The image size of the overlapping regions is 224 $\times$ 224 and the stride of the overlap sampling is 4. Once the pseudo ground truths are obtained, we randomly crop 224 $\times$ 224 regions from both the haze images and the pseudo ground truths for training our student model. We use Adam as our optimizer, with the initial learning rate set to 2e-5. We set the training step and the training batch size to 10,000 and 16, respectively. The Exponential Moving Average (EMA) weight is set to 0.9999. The thresholds $\mathbf{v}^{1}_{thr}$ and $\mathbf{v}^{2}_{thr}$ are set to 0.005 and 0.
@@
Figure 6: Qualitative results from NightEnhance’23 [17], NightDeFog’20 [53], DiT’23 [36], and our method, on the real-world dataset. Our method not only handles glow effects but also reveals detailed textures in low-light regions, such as walls and streets. Overall, results from our method look more realistic. Zoom in for better visualisation.
4.2 Quantitative Evaluation on RealNightHaze
We compare our NightHaze method on RealNightHaze with state-of-the-art (SOTA) methods, including NightEnhance [17], NightDeFog [53], NightVDM [32], Uformer [50], Restormer [58], WeatherDiffusion [35] and DiT [36]. NightEnhance [17], NightDeFog [53] and NightVDM [32] are nighttime image dehazing methods, while the other methods are image restoration backbones. WeatherDiffusion [35] and DiT [36] are diffusion-based methods. Table 1 shows that our method achieves a significant performance improvement in all non-reference metrics. For example, the MUSIQ and TRES of NightDeFog are 45.12 and 57.64, respectively. Our method achieves 52.87 of MUSIQ and 76.88 of TRES, which outperforms NightDeFog by 7.11 and 19.24, respectively. Furthermore, it can be observed that our image quality scores are closer to those of clean night images. This is because we propose self-prior learning, which brings MAE-like strong representations to nighttime image enhancement. By utilizing severe augmentation, we can generate strong network priors that demonstrate robustness against real-world haze effects, thereby leading to superior performance.
Additionally, we compare our method with other powerful image restoration backbones, including Uformer [50], Restormer [58], WeatherDiffusion [35] and DiT [36]. The experimental results are shown in Table 1. It can be observed that existing image restoration methods trained using synthetic datasets struggle to restore clear images. The main reason is that their learning strategies, based on synthetic datasets, struggle to develop strong priors. This results in inferior performance.
Table 2: Ablation studies on the RealNightHaze dataset. Five non-reference metrics, including MUSIQ [19], ManIQA [54], ClipIQA [48], HyperIQA [44], and TRES [9], and the image contrast are used for evaluation. For five non-reference metrics, higher is better. In terms of image contrast, the better it is when the contrast closely matches the statistical value of a clear image. “Ours(xx%)” means xxof the training data was severely augmented, while the remaining data was non-severely augmented. “syn-training” means training our backbone on synthetic datasets.
@@ 4.2 Quantitative Evaluation on RealNightHaze
Figure 7: Ablation studies on real-world haze images. ‘PL’ and ‘SR’ stand for self-prior learning and self-refinement, respectively. (a) Input (b) Results obtained using our self-prior learning. (c) Results obtained using our self-prior learning and self-refinement. Zoom in for better visualisation.
4.3 Qualitative Evaluation on RealNightHaze
@@
Figure 9 presents a qualitative comparison on the RealNightHaze dataset. It is evident that existing nighttime image dehazing methods cannot effectively recover a clean background in challenging areas, such as those with low-light and glow effects. The main reason is that there is a large domain gap between their synthetic datasets and nighttime haze degradation. As a result, these methods struggle to reveal a clear background in such challenging areas. In contrast, our NightHaze method not only removes nighttime haze effects but also enhances the visibility of images. This is because our self-prior learning effectively forces our network to learn strong priors. Similar to MAE, our self-prior learning is also built on augmentation. In our case, we utilize two key challenging aspects of nighttime images as augmentation factors: light effects and noise. Following the same principle as MAE, we define and employ severe augmentation to obtain a strong prior, which significantly enhances performance. Furthermore, we employ a self-refinement module to further refine our model through learning from real haze images. To be more specific, our teacher model generates high-confidence predictions from real haze images to guide the student model’s learning. This method enables our model to gradually adapt to real-world high-confidence predictions, thereby addressing more challenging scenarios.
4.4 Ablation Studies
@@
To verify the effectiveness of our self-prior learning and self-refinement, we conduct ablation studies on the RealNightHaze dataset. The experimental results are shown in Figure 7 and Table 2.
Analysis of Self-Prior Learning
@@
To verify the effectiveness of our self-prior learning, we use different ratios to mix severe and non-severe augmentation. The experimental results are shown in Table 2. It can be observed that our backbone with non-severe augmentation cannot sufficiently enhance the visibility of haze images. The MUSIQ and TRES scores are 44.67 and 62.18, outperforming the original haze images’ scores by 2.55 and 4.24. In contrast, our backbone with severe augmentation achieves significant performance improvement. The main reason for this is that severe augmentation yields strong priors that are robust to real-world haze.
Analysis of Self-Refinement
@@
We conduct ablation studies to verify the effectiveness of our self-refinement. The experimental results are shown in Figure 7. Although our self-prior learning can effectively enhance the visibility of night haze images, it still leaves some artifacts caused by over-suppression. For example, the light sources in Figure 7(b) are over-suppressed. Figure 7(c) demonstrates that our self-refinement further refines our outputs.
Figure 8: (a) Quantitative results from NightDeFog’20 [53], NightEnhance’23 [17], NightVDM’22 [32], Restormer’22 [58], DiT’23 [36] and our method, on the real-world dataset, are presented. In this evaluation, we use only the results from our self-prior learning. We utilize five non-reference image quality assessment metrics: MUSIQ [19], ManIQA [54], ClipIQA [48], HyperIQA [44], and TRES [9]. ‘Clean’ and ‘Haze’ denote statistics derived from clear night images and haze images, respectively. For all metrics, a larger value indicates better performance. (b) Quantitative results from models trained on synthetic data (blue lines) and our method with different ratios of severely augmented samples in training are presented. “Ours(xx%)” means xxof the training data was severely augmented, while the remaining data was non-severely augmented. “syn-training” means training models on synthetic datasets.
5 Discussion
In this section, we further discuss the differences between our self-prior learning and supervised learning based on synthetic data
Discussion on Training Data
Synthesizing haze requires accurate depth estimation, which is still an open challenge in real-world nighttime images. Although game engines can provide accurate depth for scenes, there remains a significant gap between these game scenes and real-world scenes. As shown in Figure 8, NF-IQA scores of results obtained from models trained on synthetic datasets have significant gaps compared to NF-IQA scores of clear night images.
In contrast, we incorporate the MAE-like framework into nighttime image enhancement and propose self-prior learning to obtain strong priors that are robust to real-world nighttime haze effects. Our main novelty is the design of severe augmentation, allowing our prior learning to become more effective. Once these robust priors are acquired, high-quality dehazing can be achieved. In our self-prior learning, we utilize two key challenging factors of nighttime images as augmentation: light effects and noise. As introduced in Section 3.1, these augmentations do not rely on any external estimation that might be inaccurate. As demonstrated in MAE [13], as long as the augmentation is severe, the priors become stronger. As shown in Figure 8, our self-prior learning significantly outperforms models trained with synthetic data.
Discussion on Degradations
Unlike synthetic data, which randomly adds degradations such as noise, our self-prior learning emphasizes the importance of augmentation severity during training. By increasing noise values to approximate the pixel intensity values of glow and light effect blended images, our learning process yields strong priors. As demonstrated in Table 2, using severe augmentation achieves significant performance improvement. Our NightHaze achieves an outstanding MUSIQ score of 52.87, significantly surpassing methods trained on synthetic datasets by a substantial margin of 19.5% in MUSIQ.
6 Conclusion
@@
This paper brings the MAE-like framework to nighttime image restoration and shows that severe augmentation during training yields strong priors that are robust to real-world night haze effects. We present NightHaze, a novel nighttime dehaizng method with self-prior learning. Our main novelty is the design of severe augmentation, which enables our prior learning to become more effective. Our self-prior learning is simple: we intentionally augment clear nighttime images by blending them with various light effects including glow and adding noise. Specifically, we blend clear images with various light maps containing artificial glow and add noise to these blended images. Inspired by MAE, which indicates that the severity of augmentation is key to obtaining strong priors, we define severe augmentation using our night-specific augmentations. We show that severe augmentation in training yields strong priors that are robust to real-world haze. Although our self-prior learning can effectively handle haze effects, it still leaves some artifacts caused by over-suppression. Our self-refinement is proposed to address this issue. It is built on the semi-supervised teacher-student framework. Unlike traditional teacher-student frameworks, we incorporate non-reference metrics to monitor each update of our student model. This approach allows us to prevent inaccurate knowledge transfer from the student to the teacher. Extensive experimental results demonstrate that our NightHaze achieves a significant performance improvement on real-world nighttime hazy datasets.
Supplementary Material
[Uncaptioned image][Uncaptioned image]
(a) Input                                   (b) Restormer [58]
[Uncaptioned image][Uncaptioned image]
(c) DiT [36]                                (d) NightDeFog [53]
[Uncaptioned image][Uncaptioned image]
(e) NightEnhance [17]                             (f) Ours
Figure 9: Qualitative results from NightEnhance’23 [17], NightDeFog’20 [53], Restormer’22 [58], DiT’23 [36], and our method, on the real-world dataset. Our method not only handles glow effects but also reveals detailed textures in low-light regions. Overall, results from our method look more realistic. Zoom in for better visualisation.
More Results on Real-world Datasets
In this section, we compare our NightHaze method on the RealNightHaze dataset with state-of-the-art (SOTA) methods, including NightEnhance [17], NightDeFog [53], Restormer [58], and DiT [36]. NightEnhance [17] and NightDeFog [53] are nighttime image dehazing methods, while the other methods are image restoration backbones. The experimental results are shown in Figs. 9, 11 and 10. It can be observed that existing nighttime image dehazing methods cannot effectively recover a clean background in challenging areas, such as those with low-light and glow effects. In contrast, our NightHaze method not only removes nighttime haze effects but also enhances the visibility of images.
Figure 10: Qualitative results from NightEnhance’23 [17], NightDeFog’20 [53], Restormer’22 [58], DiT’23 [36], and our method, on the real-world dataset. Our method not only handles glow effects but also reveals detailed textures in low-light regions. Overall, results from our method look more realistic. Zoom in for better visualisation.
Figure 11: Qualitative results from NightEnhance’23 [17], NightDeFog’20 [53], Restormer’22 [58], DiT’23 [36], and our method, on the real-world dataset. Our method not only handles glow effects but also reveals detailed textures in low-light regions. Overall, results from our method look more realistic. Zoom in for better visualisation.
(a) Input
Refer to captionRefer to caption
(b) UNIE [18]
Refer to captionRefer to caption
(c) Ours
Figure 12: Qualitative results from UNIE [18] and our method, on the real-world dataset. Overall, results from our method look more realistic. Zoom in for better visualisation.
Extension to Other Tasks
To verify the applicability of our method to other tasks, we also conduct experiments on backlit image enhancement [29] and night image enhancement [18].
Night Image Enhancement
The goal of night image enhancement is to suppress light effects in bright regions while, simultaneously, boosting the intensity in dark regions. We use the LightEffects [18] dataset to evaluate the performance. In this experiment, we use only the results from our self-prior learning. The experimental results are presented in Figs. 12, 13, 14 and 15. We can observe that our method can effectively enhance the visibility of night images.
(a) Input
Refer to captionRefer to caption
(b) UNIE [18]
Refer to captionRefer to caption
(c) Ours
Figure 13: Qualitative results from UNIE [18] and our method, on the real-world dataset. Overall, results from our method look more realistic. Zoom in for better visualisation.
(a) Input
Refer to captionRefer to caption
(b) UNIE [18]
Refer to captionRefer to caption
(c) Ours
Figure 14: Qualitative results from UNIE [18] and our method, on the real-world dataset. Overall, results from our method look more realistic. Zoom in for better visualisation.
(a) Input
Refer to captionRefer to caption
(b) UNIE [18]
Refer to captionRefer to caption
(c) Ours
Figure 15: Qualitative results from UNIE [18] and our method, on the real-world dataset. Overall, results from our method look more realistic. Zoom in for better visualisation.
(a) Input
Refer to captionRefer to caption
(b) CLIP-LiT [29]
Refer to captionRefer to caption
(c) Ours
Figure 16: Qualitative results from CLIP-LiT [29] and our method, on the real-world dataset. Overall, results from our method look more realistic. Zoom in for better visualisation.
(a) Input
Refer to captionRefer to caption
(b) CLIP-LiT [29]
Refer to captionRefer to caption
(c) Ours
Figure 17: Qualitative results from CLIP-LiT [29] and our method, on the real-world dataset. Overall, results from our method look more realistic. Zoom in for better visualisation.
Backlit Image Enhancement
For backlit image enhancement, We evaluate our method using the Backlit300 dataset [29], which consists of 305 backlit images. Following the evaluation protocol from [29], we use a non-reference metric, MUSIQ [20], to quantitatively compare our NightHaze method with state-of-the-art (SOTA) methods. In this evaluation, we use only the results from our self-prior learning. The experimental results are presented in Table 3, while Figs. 16 and 17 illustrate the qualitative results. It can be observed that our method can effectively enhance the visibility of backlit images.
Table 3: Quantitative comparison on the Backlit300 test dataset [29]. Higher is better.
Method Input Afifi et al. [1] MIT5K [64] URetinex-Net [52] Zero-DCE [12] Zero-DCE++ [25]
MUSIQ 51.900 51.930 50.354 51.551 51.250 48.435
SCI-easy [33] SCI-medium [33] SCI-diffucult [33] EnlightenGAN [16] ExCNet [63] CLIP-LIT [29] Ours
50.642 43.493 49.428 48.308 50.278 52.921 54.470

### v2

@@ Introduction
Although our self-prior learning effectively handles haze effects, it occasionally results in artifacts, particularly in the form of over-suppression. To address this issue, we propose a teacher-student framework (Lin et al. 2024; Chen et al. 2025) to further refine our network using real haze images. The network parameter from our self-prior learning is used to initialize a teacher model $w_{T}$ and a student model $w_{S}$. The purpose of the teacher model is to generate the pseudo ground truth based on confidence score, while the student model is trained on high-confidence predictions and their corresponding inputs. By gradually learning from high-confidence predictions, our network is allowed to address more challenging regions.
Given a real-world hazy dataset $\mathbf{D}_{\rm uh}=\{\mathbf{x}^{\rm uh}_{i}\}^{N_{\rm uh}}_{i=1}$, where $\mathbf{x}^{\rm uh}_{i}$ is the $i$-th haze image and $N_{\rm uh}$ is the number of haze images, we utilize the teacher model $w_{\rm T}$ to generate predictions $\mathbf{y}^{\rm uh}_{i}$ and confidence maps $\mathbf{u}^{\rm uh}_{i}$. Specifically, we first sample overlapping regions from $\mathbf{x}^{\rm uh}_{i}$ and input the set of overlapping regions into our teacher model $w_{T}$ to generate predictions. Then, we calculate the mean values and variance of overlapping regions as predictions and confidence maps. Subsequently, we use a threshold $\mathbf{v}^{1}_{thr}$ to convert the confidence map $\mathbf{u}^{\rm uh}_{i}$ to a binary mask $\mathbf{m}^{\rm uh}_{i}$.
During the training stage, we use the noise and haze augmentation to degrade the haze images $\mathbf{x}^{\rm uh}_{i}$. The degraded haze images $\mathbf{\hat{x}}^{\rm uh}_{i}$ are then fed into our student model for training. The loss function can be formulated as:
where $N$ is the number of samples in the training batch and $\mathbf{\hat{y}}^{\rm uh}_{i}$ is the output of the degraded haze images $\mathbf{\hat{x}}^{\rm uh}_{i}$. Once the losses are obtained, traditional teacher-student frameworks use back-propagation to update the student model $w_{S}$. Then they use the Exponential Moving Average (EMA) to update the teacher model. However, the student model is trained using augmented real-world haze inputs. Since the augmentation is random, the updates of the student model in some training steps may be wrong and thus affect the performance of the teacher model by EMA. To avoid this, we introduce a non-reference metric to monitor each update of the student model, which can be formulated as:
Experiments
Due to the lack of ground truth, previous methods typically use synthetic datasets for quantitative evaluation and real-world datasets for qualitative evaluation. However, the quantitative evaluation on synthetic datasets cannot effectively represent the real-world dehazing capabilities. Motivated by this, we introduce a new benchmark to quantitatively evaluate the dehazing performance on real-world nighttime haze datasets. Specifically, we use five non-reference image quality assessments, including MUSIQ (Ke et al. 2021), ManIQA (Yang et al. 2022), ClipIQA (Wang, Chan, and Loy 2023), HyperIQA (Su et al. 2020), and TRES (Golestaneh, Dadsetan, and Kitani 2022), along with image contrast, to evaluate the performance of existing methods.
Our real-world nighttime haze datasets, RealNightHaze, is collected from internet and existing nighttime dehazing methods (Zhang et al. 2020; Jin et al. 2023), and consists of 440 real-world nighttime haze images.
(a) Inputs
(b) DiT
(c) NightDeFog
(d) NightEnhance
(e) Ours
@@
Figure 6: Qualitative results from NightEnhance’23 (Jin et al. 2023), NightDeFog’20 (Yan, Tan, and Dai 2020), DiT’23 (Peebles and Xie 2023), and our method, on the real-world dataset. Our method not only handles glow effects but also reveals detailed textures in low-light regions, such as walls and streets. Zoom in for better visualisation.
Implementation Details
Our NightHaze includes two parts: self-prior learning and self-refinement. We adopt the MAE network (He et al. 2022) as our backbone. In our self-prior learning, the size of input images is 224 $\times$ 224. During the training stage, we use Adam as our optimizer and set the initial learning rate to 1.5e-4. The total training steps and the training batch size are set to 20,000 and 128, respectively. Our self-prior learning consists of three parts: the light map, the blending weight map, and noise.
Light Map At each training step, we randomly select an atmospheric light map for every clear image. Then, we select various regions and amplify their brightness using Gaussian kernels. The number of selected regions is between 2 and 10, and the size of Gaussian kernels ranges from 15 to 80.
Blending Weight Map We initially select a blending weight $t$ from a uniform distribution ($T_{\rm low}$, $T_{\rm high}$) and expand it into a map. Then, we randomly adjust the blending weights of several regions. $T_{\rm low}$ and $T_{\rm high}$ are set to 0.001 and 0.1, respectively. We set the number of the selected regions to 8. The size of each region is 64 $\times$ 64 and the range of the changed value is a uniform distribution (0, 0.04).
Noise During the training stage, we use Gaussian noise to augment our input images. We set $W_{n}$ to 0.1. Thus, the value range of our $\epsilon$ becomes to (0, 0.03).
Self-Refinement For a haze image of size 256 $\times$ 256, we first sample overlapping regions within the input haze image. The image size of the overlapping regions is 224 $\times$ 224 and the stride of the overlap sampling is 4. Once the pseudo ground truths are obtained, we randomly crop 224 $\times$ 224 regions from both the haze images and the pseudo ground truths for training our student model. We use Adam as our optimizer, with the initial learning rate set to 2e-5. We set the training step and the training batch size to 10,000 and 16, respectively. The EMA weight is set to 0.9999. The thresholds $\mathbf{v}^{1}_{thr}$ and $\mathbf{v}^{2}_{thr}$ are set to 0.005 and 0.
@@
Table 2: Ablation studies on the RealNightHaze dataset. For non-reference metrics, higher is better. In terms of image contrast, the better it is when the contrast closely matches the statistical value of a clear image. “xxsevere” means xxof the training data was severely augmented, while the remaining data was non-severely augmented. “syn-training” means training our backbone on synthetic datasets.
Quantitative Evaluation on RealNightHaze
We compare our NightHaze method on RealNightHaze with state-of-the-art (SOTA) methods, including NightEnhance (Jin et al. 2023), NightDeFog (Yan, Tan, and Dai 2020), NightVDM (Liu et al. 2022), Uformer (Wang et al. 2022), Restormer (Zamir et al. 2022), WeatherDiffusion (Özdenizci and Legenstein 2023) and DiT (Peebles and Xie 2023). Table 1 shows that our method achieves a significant performance improvement in all non-reference metrics. For example, the MUSIQ and TRES of NightDeFog are 45.76 and 65.83, respectively. Our method achieves 52.87 of MUSIQ and 76.88 of TRES, which outperforms NightDeFog by 7.11 and 11.05, respectively. Furthermore, it can be observed that our image quality scores are closer to those of clean night images. This is because we propose self-prior learning, which brings MAE-like strong representations to nighttime image enhancement. By utilizing severe augmentation, we can generate strong priors that demonstrate robustness against real-world haze effects, thereby leading to superior performance.
Additionally, we compare our method with other powerful image restoration backbones, including Uformer (Wang et al. 2022), Restormer (Zamir et al. 2022), WeatherDiffusion (Özdenizci and Legenstein 2023) and DiT (Peebles and Xie 2023). The experimental results are shown in Table 1. It can be observed that existing image restoration methods trained using synthetic datasets struggle to restore haze images. The main reason is that their learning strategies, based on synthetic datasets, struggle to develop strong priors. This results in inferior performance.
Qualitative Evaluation on RealNightHaze
@@
Figure 6 presents a qualitative comparison on RealNightHaze. It is evident that existing nighttime image dehazing methods cannot effectively recover a clean background in challenging areas, such as those with low-light and glow effects. The main reason is that there is a large domain gap between their synthetic datasets and nighttime haze degradation. In contrast, our NightHaze method not only removes nighttime haze effects but also enhances the visibility of images. This is because our self-prior learning effectively forces our network to learn strong priors. Similar to MAE, our self-prior learning is also built on augmentation. In our case, we utilize two key challenging aspects of nighttime images as augmentation factors: light effects and noise. Following the same principle as MAE, we define and employ severe augmentation to obtain a strong prior, which significantly enhances performance. Furthermore, we employ a self-refinement module to further refine our model through learning from real haze images. To be more specific, our teacher model generates high-confidence predictions from real haze images to guide the student model’s learning. This method enables our model to gradually adapt to real-world high-confidence predictions, thereby addressing more challenging scenarios.
(a) Inputs
(b) PL
(c) PL+SR
Figure 7: Ablation studies on real-world haze images. ‘PL’ and ‘SR’ are self-prior learning and self-refinement, respectively. (a) Input (b) Results obtained using our self-prior learning. (c) Results obtained using our self-prior learning and self-refinement. Zoom in for better visualisation.
Ablation Studies
@@
To verify the effectiveness of our key ideas, we conduct ablation studies on the RealNightHaze dataset. The experimental results are shown in Figure 7 and Table 2.
@@
Analysis of Self-Prior Learning To verify the effectiveness of our self-prior learning, we use different ratios to mix severe and non-severe augmentation. The experimental results are shown in Table 2. It can be observed that our backbone with non-severe augmentation cannot sufficiently enhance the visibility of haze images. The MUSIQ and TRES scores are 44.67 and 62.18, outperforming the original haze images’ scores by 2.55 and 4.24. In contrast, our backbone with severe augmentation achieves significant performance improvement. The main reason for this is that severe augmentation yields strong priors that are robust to real haze.
@@
Analysis of Self-Refinement We conduct ablation studies to verify the effectiveness of our self-refinement. The experimental results are shown in Figure 7. Although our self-prior learning can effectively enhance the visibility of night haze images, it still leaves some artifacts caused by over-suppression. For example, the light sources in Figure 7(b) are over-suppressed. Figure 7(c) demonstrates that our self-refinement further refines our outputs.
Conclusion
@@
This paper brings the MAE-like framework to nighttime image restoration and shows that severe augmentation during training yields strong priors that are robust to real-world night haze effects. We present NightHaze, a novel nighttime dehaizng method with self-prior learning. Our self-prior learning is simple: we intentionally augment clear nighttime images by blending them with various light effects including glow and adding noise. Specifically, we blend clear images with various light maps containing artificial glow and add noise to these blended images. Inspired by MAE, which indicates that the severity of augmentation is key to obtaining strong priors, we define severe augmentation using our night-specific augmentations. We show that severe augmentation in training yields strong priors that are robust to real-world haze. Although our self-prior learning can effectively handle haze effects, it still leaves some artifacts caused by over-suppression. Our self-refinement is proposed to address this issue. It is built on the semi-supervised teacher-student framework. Unlike traditional teacher-student frameworks, we incorporate non-reference metrics to monitor each update of our student model. This approach allows us to prevent inaccurate knowledge transfer from the student to the teacher. Extensive experimental results demonstrate that our NightHaze achieves a significant performance improvement on real-world nighttime hazy datasets.